OpenStudy (anonymous):

Hey could you please tell how me to find the magnitude and phase of H(jw) = j10w/(j10w +1) thank you x

OpenStudy (phi):

I assume j= sqrt(-1), and w is omega. put the expression into rectangular form (multiply top and bottom by the complex congugate of (1+10w j) then find the mag of a+j b = sqrt(a^2+b^2) and phase = atan(b/a)